Electric Force of Three Collinear Points Ranking Task Part A Rank ธ ฝ0sebitations et uleenk darges on the basis of th..lectric tore acin9 or, g, Defne sieces paleving to the aghr us posb. and forces pointing to the left as Rank in increasing order by placing the most negative left and the mos4 Positive on the light. โฉ lank items as equivalent overlap them we
Answer #1

Force applied by the q2 on q1 is F21

Force applied by the q3 on q1 is F31

Now let the distance between the q1 and q2 is d

Then the distance between the q1 and q3 = 2d

As all charges are identical i.e q1=q2=q3

The force F21 is

F21= q1*q2*k / d^2

Wheree k=1/4π\varepsilon

Now the force F31 can be written as

F31= q1*q3*k/(2d)^2= q1*q2*k/4d^2

F31= F21/4

Hence F21 > F31

F21 = force applied by q2 on q1 is at 1st rank

F31= force applied by q3 on q1 is at 2nd rank